Cu2S + O2à2Cu + SO2

If the reaction of .760 Kg of copper sulfide(I) with excess oxygen produces 0.310 kg of copper metal, what is the percentage yield?

  1. 40.8%
  2. 39.9%
  3. 204%
  4. 51.1%
  5. 102%
Answer #1

Answer

51.1%

Explanation

Cu2S + O2 -------> 2Cu + SO2

Stochiometrically, 1mole of Cu2S gives 2moles of Cu

given moles of Cu2O = 760g/159.16= 4.775mol

moles of Cu can be maximum obtained by 4.775moles of Cu2O = 2 × 4.775mol = 9.55mol

mass of Cu can be maximum obtained = 9.55mol × 63.55g/mol = 606.9g = 0.6069kg

Theoretical yield = 0.6069kg

percentage yield = ( 0.310kg/0.6069kg) × 100 = 51.1%